OpenStudy (anonymous):

For what values of a and b are y = a bx and y = 2 + x tangent at x = 0.

OpenStudy (anonymous):

it should be y=a*b^x

OpenStudy (anonymous):

please clarify you are asking Y=a*b^x to be tangent on Y=2+x at x=0.. OR both Y=a*b^x & Y=2+x both be a tangent on x=0

OpenStudy (anonymous):

y=a*b^x is tangent at y=2+x at x=0

OpenStudy (anonymous):

i need to find the values for a and b

OpenStudy (anonymous):

i know that the point is (0,4) and the derivative of of y=ab^x is y'=ab^x*ln(b)

OpenStudy (anonymous):

so the slope at x=0 is 1

OpenStudy (anonymous):

here derivative of y=2+x will give the slope of y=a*b^y

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

why would you need the derivative of y=2+x

OpenStudy (anonymous):

its just a line

OpenStudy (anonymous):

see on derivating a function we get the slope of its tangent at a given point.

OpenStudy (anonymous):

yeah but y=2+x is the line tangent to ab^x

OpenStudy (anonymous):

according to you question "y=a*b^x is tangent at y=2+x at x=0 "this would be the approach.

OpenStudy (anonymous):

no dude you dont know

OpenStudy (anonymous):

you dont need the derivative of y=2+x we need the derivative of y=ab^x

OpenStudy (anonymous):

yeah that i knw but in question u typed something else..anyways i will get back to u in 5 minutes ..need to solve on paper.

OpenStudy (anonymous):

i did not!

OpenStudy (anonymous):

okay.but plz clarify one more thing the curve is Y=a*b^x or y=(a*b)^x.?

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

dont wory about

OpenStudy (anonymous):

my answer.a=2;b=e^0.5...is it correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanx.